OpenStudy (anonymous):

solve using any algebraic method x+3/4+y-1/3=1 and 2x-y=12 is the equation, please explain thanks

OpenStudy (agreene):

\[x+3/4+y-1/3=1\] \[x+y+5/12=1\] \[x=7/12-y\]plug this in to the other eqn. \[2x-y=12\] \[2(7/12-y)-y=12\] \[7/6-2y-y=12\] \[7/6-3y=12\] \[-3y=65/6\] \[y=-65/18\] plug this y=-65/18 in to the x= equation \[x=7/12-(-65/18)\] \[x=151/36\] x=151/36 or 4.19444 y= -65/18 or -3.6111

OpenStudy (anonymous):

uhmm, how did you get x+y+5/12=1

OpenStudy (agreene):

3/4-1/3 = 5/12